# On Supports for Graphs of Bounded Genus [^1] Parts of this work appeared in “On Hypergraph Supports” available at https://arxiv.org/abs/2303.16515, which has now been split into two papers. This is the first paper in the series, and the second paper is available at https://arxiv.org/abs/2504.05039v2

### v2

@@ 7.1 Algorithms for Packing and Covering Problems
Pick a largest sub-collection $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ s.t. |$|\{S\in\mathcal{S}^{\prime}:S\ni x\}|\leq 1$ for all $x\in X$.
@@
In abstract set systems, there is essentially no difference between the dual pairs of problems Point Packing/Set Packing, or Hitting Set/Set Cover. The Set Packing problem contains as a special case, the Independent Set problem on graphs, and is therefore hard to approximate beyond $n^{1-\epsilon}$ for any $\epsilon>0$ unless NP$=$ZPP [hastad1999clique], and the Set Cover problem admits an $O(\log n)$-approximation [chvatal1979greedy, lov1975notdef], and that this is tight [feige1998threshold].
@@ 7.1.1 PTAS via local-search
Let $\mathcal{L}\subseteq\mathcal{H}$ be a solution returned by the local-search algorithm and let $\mathcal{O}\subseteq\mathcal{H}$ be an optimal solution. Since we are interested in bounding $|\mathcal{L}|$ by $|\mathcal{O}|$, we remove the subgraphs in $\mathcal{L}\cap\mathcal{O}$ and replace the capacity of each $K\in\mathcal{K}$ by its residual capacity: $\textrm{cap}^{\prime}(K)=\textrm{cap}(K)-|\{X:X\cap K\neq\emptyset\mbox{ and }X\in\mathcal{L}\cap\mathcal{O}\}|$. We also remove the subgraphs in $\mathcal{K}$ whose residual capacity is 00. Let $\mathcal{L}^{\prime}=\mathcal{L}\setminus(\mathcal{L}\cap\mathcal{O})$ and $\mathcal{O}^{\prime}=\mathcal{O}\setminus(\mathcal{L}\cap\mathcal{O})$. If we prove $|\mathcal{L}^{\prime}|\leq(1+\epsilon)|\mathcal{O}^{\prime}|$, then it implies that $|\mathcal{L}|\leq(1+\epsilon)|\mathcal{O}|$. Hence, we assume below that $\mathcal{L}\cap\mathcal{O}=\emptyset$, and each $K\in\mathcal{K}$ has positive capacity bounded above by $\Delta$.
@@
The graph $\Xi$ may not have bounded genus. Nevertheless, we show that $\Xi$ has a sub-linear separator. Each subgraph $H\in\mathcal{L}\cup\mathcal{O}$ assigns a weight of $1/|H|$ to the vertices of $H$ in the intersection support $\tilde{Q}$. The weight of a vertex in $\tilde{Q}$ is the sum of the contributions it receives from the subgraphs containing it. Let $w(K)$ denote the weight of a vertex $K\in\tilde{Q}$ and let $W=|\mathcal{L}\cup\mathcal{O}|$ denote the total weight of the vertices in $\tilde{Q}$. Since $\tilde{Q}$ has bounded genus, by Theorem 3, it has a weighted separator of size $O(\sqrt{g|\mathcal{K}^{\prime}|})$. That is, a set $S\subseteq V(\tilde{Q})$ of size $O(\sqrt{g|\mathcal{K}^{\prime}|})$ s.t. $V(\tilde{Q})\setminus S$ separates $\tilde{Q}$ into disconnected components that can be partitioned into two groups $A$ and $B$ with no edges between them, and that $w(A),w(B)\leq\alpha W$, for some constant $10<\alpha<$ 1.
@@
For the Independent Set and Dominating Set problems, the algorithm follows directly via local-search. Notably, for the Dominating Set problem, we require an intersection support. For the Vertex Cover problem, we use the well-known half-integrality [nemhauser1975vertex] of the standard LP-relaxation for Vertex Cover. That is, there is an optimal solution to the LP $\min\{\sum x_{v}:x_{u}+x_{v}\geq 1\;\forall\{u,v\}\in E,x_{u}\geq 0\}$ where the $x_{u}$ take values in $\{0,1/2,1\}$. The vertices whose values are either $1$ or $1/2$ is a feasible vertex cover. We remove the vertices of weight 00, take all vertices of weight 1, and in the graph induced on the vertices of weight $1/2$, it follows that any vertex cover requires at least $1/2$ the vertices. Therefore, a $(1-\epsilon)$-approximation to the Independent Set problem on this induced subgraph implies a $(1+\epsilon)$-approximation to the Vertex Cover problem. These ideas have been observed previously by Bar-Yahuda et al., [bar2011minimum] to obtain a $3/2$-approximation for the Vertex Cover problem on the intersection graph of axis-aligned rectangles.
@@ 7.2 Regions on an Oriented Surface
A set of bounded regions $\mathcal{D}$ is a set of pseudodisks if their boundaries are pseudocircles. Note that each $D\in\mathcal{D}$ defines a simply connected region [^8] A region $R$ is simply connected if any simple loop can be continuously contracted to a point so that the loop lies in $R$ throughout the contraction process. Intuitively, a simply connected region does not have holes. such that for any $D,D^{\prime}\in\mathcal{D}$, $D\setminus D^{\prime}$ and $D^{\prime}\setminus D$ are connected.
If we only insist on the last property, we obtain a set of non-piercing regions, a generalization of pseudodisks. Recall Definition 17 that a set $\mathcal{D}$ of bounded, connected regions in the plane are a set of non-piercing regions if for any $D,D^{\prime}\in\mathcal{D}$, $D\setminus D^{\prime}$ and $D^{\prime}\setminus D$ are connected. However, we impose the restriction that the boundaries of $D$ and $D^{\prime}$ intersect a finite number of times for any $D,D^{\prime}\in\mathcal{D}$. Note that the sets in $\mathcal{D}$ need not be simply connected. Figure 5 shows examples of pseudodisks, piercing regions, and non-piercing regions in the plane. Pseudodisks and non-piercing regions generalize several well-studied geometric objects in the plane such as disks, squares, homothets of a convex set, etc.
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be two collections of regions on an surface. The intersection hypergraph of the regions is a hypergraph (,$(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$, i.e., it takes a vertex for each $H\in\mathcal{H}$, and a hyperedge $\mathcal{H}_{K}$ for each $K\in\mathcal{K}$ where $\mathcal{H}_{K}=\{H\in\mathcal{H}:H\cap K\neq\emptyset\}$.
@@
Figure 6: The subgraphs $A$ colored orange, $B$ colored green, and $C$ colored blue are touching. Observe that $A\setminus B$ is connected, while $B\setminus A$ induces two connected components. There is an input point on the boundary of the three regions. The regions are weakly non-piercing, but are not non-piercing.
@@ 8 Conclusion
There are several intriguing open questions and research directions, and we mention a few: We do not know if the algorithm to construct a dual or intersection support in the bounded-genus case runs in polynomial time.

### v3

@@ 7.1 Algorithms for Packing and Covering Problems
Pick a largest sub-collection $\mathcal{S}^{\prime}\subseteq\mathcal{S}$ s.t. $|\{S\in\mathcal{S}^{\prime}:S\ni x\}|\leq 1$ for all $x\in X$.
@@
In abstract set systems, there is essentially no difference between the dual pairs of problems Point Packing/Set Packing, or Hitting Set/Set Cover. The Set Packing problem contains as a special case, the Independent Set problem on graphs, and is therefore hard to approximate beyond $n^{1-\epsilon}$ for any $\epsilon>0$ unless NP==ZPP [hastad1999clique], and the Set Cover problem admits an $O(\log n)$-approximation [chvatal1979greedy, lov1975notdef], and that this is tight [feige1998threshold].
@@ 7.1.1 PTAS via local-search
Let $\mathcal{L}\subseteq\mathcal{H}$ be a solution returned by the local-search algorithm and let $\mathcal{O}\subseteq\mathcal{H}$ be an optimal solution. Since we are interested in bounding $|\mathcal{L}|$ by $|\mathcal{O}|$, we remove the subgraphs in $\mathcal{L}\cap\mathcal{O}$ and replace the capacity of each $K\in\mathcal{K}$ by its residual capacity: $\textrm{cap}^{\prime}(K)=\textrm{cap}(K)-|\{X:X\cap K\neq\emptyset\mbox{ and }X\in\mathcal{L}\cap\mathcal{O}\}|$. We also remove the subgraphs in $\mathcal{K}$ whose residual capacity is 0. Let $\mathcal{L}^{\prime}=\mathcal{L}\setminus(\mathcal{L}\cap\mathcal{O})$ and $\mathcal{O}^{\prime}=\mathcal{O}\setminus(\mathcal{L}\cap\mathcal{O})$. If we prove $|\mathcal{L}^{\prime}|\leq(1+\epsilon)|\mathcal{O}^{\prime}|$, then it implies that $|\mathcal{L}|\leq(1+\epsilon)|\mathcal{O}|$. Hence, we assume below that $\mathcal{L}\cap\mathcal{O}=\emptyset$, and each $K\in\mathcal{K}$ has positive capacity bounded above by $\Delta$.
@@
[…]
@@
For the Independent Set and Dominating Set problems, the algorithm follows directly via local-search. Notably, for the Dominating Set problem, we require an intersection support. For the Vertex Cover problem, we use the well-known half-integrality [nemhauser1975vertex] of the standard LP-relaxation for Vertex Cover. That is, there is an optimal solution to the LP $\min\{\sum x_{v}:x_{u}+x_{v}\geq 1\;\forall\{u,v\}\in E,x_{u}\geq 0\}$ where the $x_{u}$ take values in $\{0,1/2,1\}$. The vertices whose values are either $1$ or $1/2$ is a feasible vertex cover. We remove the vertices of weight 0, take all vertices of weight 1, and in the graph induced on the vertices of weight $1/2$, it follows that any vertex cover requires at least $1/2$ the vertices. Therefore, a $(1-\epsilon)$-approximation to the Independent Set problem on this induced subgraph implies a $(1+\epsilon)$-approximation to the Vertex Cover problem. These ideas have been observed previously by Bar-Yahuda et al., [bar2011minimum] to obtain a $3/2$-approximation for the Vertex Cover problem on the intersection graph of axis-aligned rectangles.
@@ 7.2 Regions on an Oriented Surface
A set of bounded regions $\mathcal{D}$ is a set of pseudodisks if their boundaries are pseudocircles. Note that each $D\in\mathcal{D}$ defines a simply connected region [^8] A region $R$ is simply connected if any simple loop can be continuously contracted to a point so that the loop lies in $R$ throughout the contraction process. Intuitively, a simply connected region does not have holes. such that for any $^{\prime}\in\mathcal{D}, DDD\setminus D^{\prime}$ and $D^{\prime}\setminus D$ are connected.
If we only insist on the last property, we obtain a set of non-piercing regions, a generalization of pseudodisks. Recall Definition 17 that a set $\mathcal{D}$ of bounded, connected regions in the plane are a set of non-piercing regions if for any $^{\prime}\in\mathcal{D}, DDD\setminus D^{\prime}$ and $D^{\prime}\setminus D$ are connected. However, we impose the restriction that the boundaries of $D$ and $D^{\prime}$ intersect a finite number of times for any $D,D^{\prime}\in\mathcal{D}$. Note that the sets in $\mathcal{D}$ need not be simply connected. Figure 5 shows examples of pseudodisks, piercing regions, and non-piercing regions in the plane. Pseudodisks and non-piercing regions generalize several well-studied geometric objects in the plane such as disks, squares, homothets of a convex set, etc.
@@
Let $\mathcal{H}$ and $\mathcal{K}$ be two collections of regions on an surface. The intersection hypergraph of the regions is a hypergraph $(\mathcal{H},\{\mathcal{H}_{K}\}_{K\in\mathcal{K}})$, i.e., it takes a vertex for each $H\in\mathcal{H}$, and a hyperedge $\mathcal{H}_{K}$ for each $K\in\mathcal{K}$ where $\mathcal{H}_{K}=\{H\in\mathcal{H}:H\cap K\neq\emptyset\}$.
@@
Figure 6: The subgraphs AA colored orange, $B$ colored green, and $C$ colored blue are touching. Observe that $A\setminus B$ is connected, while $B\setminus A$ induces two connected components. There is an input point on the boundary of the three regions. The regions are weakly non-piercing, but are not non-piercing.
@@ 8 Conclusion
There are several intriguing open questions and research directions, and we mention a few: We do not know if the algorithms to construct a primal, dual or intersection support run in polynomial time.